OpenStudy (anonymous):

Find the nth degree polynomial with real coefficients satisfying the given conditions. n=3, 4 and 2i are zeros, f(-1)=50 The expanded and simplified polynomial is f(x)=_________________

OpenStudy (antonio_xx2):

f(x) = x³ -4x²- 12x +80

OpenStudy (anonymous):

that is wrong

OpenStudy (anonymous):

The degree is 3, so you have a polynomial of the form \[ax^3+bx^2+cx+d\] You know the roots are \(x=4\) and \(x=\pm2i\) (complex roots come in conjugate pairs), so you can write it as \[a(x-4)(x-2i)(x+2i)\] You know that \(f(-1)=50\), so \[a(-1-4)(-1-2i)(-1+2i)=50~~\Rightarrow~~a=-2\] So, the polynomial (unexpanded) is \[f(x)=-2(x-4)(x-2i)(x+2i)\]